# Quantum Optimization Benchmark Library The Intractable Decathlon

### v1

@@ 5.1 Low Autocorrelation Binary Sequences
Figure 17: (LABS) Comparison of the time-to-solution for LABS instance ranging from size 10 to 30 using BF-DCQO and QAOA across different numbers of layers: $p=1$, $p=4$, $p=8$, and $p=12$. The data for QAOA is sourced from [210]. Note that the BF-DCQO times reflect idealized QPU runtime estimates without considering transpilation.
The quantum adiabatic algorithm (QAA) evolves an easy-to-prepare ground state of an initial Hamiltonian $H_{i}$ to the unknown ground state of a target Hamiltonian $H_{f}$, which should correspond to the optimal solution of an optimization problem of interest. This evolution is described by a time-dependent Hamiltonian of the form
@@
An extension of DCQO is bias-field DCQO (BF-DCQO), which is an iterative algorithm [40, 202, 117]. In BF-DCQO, the information from the final evolved state is fed back as a bias for a new initial state such that the initial Hamiltonian for another DCQO iteration is adapted accordingly. In addition to that, we consider a classical post-processing scheme where we perform a local search [55, 62, 244] on the $n_{keep}$ lowest energy bitstrings obtained from BF-DCQO at each iteration. This local search can be seen as a zero-temperature simulated annealing with $n_{s}$ sweeps.
@@
Figure 18: (LABS) Comparison of the time-to-solution for BF-DCQO, Gurobi, and CPLEX. The data for the classical optimizers was sourced from [210]. Note that the BF-DCQO times reflect idealized QPU runtime estimates without considering transpilation.
@@
Table 11: (LABS) Summary of results for ten experimental trials of BF-DCQO on the size $N=20$ LABS problem executed on ibm__\__marrakesh. The overall runtime consists of pre-processing time, which includes circuit construction and transpilation, QPU runtime, which includes the payload circuit preparation and execution time, and post-processing time, which includes the local search post-processing. Further $p_{gs}$ shows the probability to sample the best solution–here the optimal one.
[…]
@@
Table 12: (LABS) Report of benchmark submission for the LABS instance with $N=20$, also available in QOBLIB [193]. All runtimes are reported as the average over the 10 runs.
Since the Hamiltonian corresponding to LABS is dense with both 2-local and 4-local terms, its realization is challenging due to the large number of entangling gates required. Therefore, we restrict the counterdiabatic terms to include only 2-local terms. We used the 156-qubit ibm__\__marrakesh device with optimization level 2 transpilation, dynamical decoupling based on an XpXm sequence, and Qiskit 1.3.1. To facilitate the experimental implementation, we adopted the following strategy for initializing the bias fields. First, we created a random bitstring and applied a local greedy search with $n_{\text{s}}=5$. Then, we use the resulting bitstring to compute the initial bias fields. We set the initial number of shots to $n_{\text{shots}}=500$ and $n_{\text{s}}=2$. The $n_{\text{keep}}$ parameter is chosen as $50\%$ and $n_{\text{shots}}$ is increased by $500$ after each iteration. In Table 11, we present the results for ten different bias-field initializations. For the selected set of parameters, all the initializations resulted in a non-zero success probability. The $N=20$ LABS problem has an eight-fold degenerate ground state, which can be divided into two distinct independent groups based on spin-flip and reflection symmetry. In our results, we measure optimal sequences from both groups. The average overall runtime to measure the optimal solution was $61.224$ seconds. However, excluding the payload quantum circuit preparation, the average runtime becomes $9.1$ seconds, which is still better than the two exact classical methods shown in Figure 18.
Figure 19: (LABS) Quantum simulation (top) hardware (bottom) results for $n=4$ (left) and $n=5$ (right). The error markers with $10^{-12}$ error indicate an exact decomposition (i.e., zero error). The x-axis indicates the problem instance BX_sparse/dense_Z, where X is the size of the doubly stochastic matrix (i.e., $n$), sparse/dense is the number of permutations used to generate the doubly stochastic matrix ($n$ for sparse and $n^{2}$ for dense), and Z is the instance number.
@@ 5.2 Minimum Birkhoff Decomposition
In the experiments presented below, we employed Optuna 4.1 [10] as a black-box solver for Eq. (29) to optimize the parameters $\theta$. After sampling the permutation matrices, the coefficients $c_{1},\dots,c_{k}$ are computed classically in two steps. In the first step (black-box step 1), we use CPLEX 22.1.1 [47] to minimize $\|sD-\sum_{i=1}^{k}u_{i}P_{i}\|_{2}$ with respect to $u_{i}\in\{0,1,2,\dots,s\}$ subject to $0\leq u_{i}\leq s$ and $\sum_{i=1}^{k}u_{i}=s$ where $s$ is a given integer in the dataset that makes $sD$ and interger-valued matrix. In the second step (black-box step 2), we use again CPLEX 22.1.1 to find the coefficients $c_{i}\in[0,1]$ that provide the smallest decomposition subject to $\|D-\sum_{i=1}^{k}c_{i}P_{i}\|_{2}\leq\|D-\frac{1}{s}\sum_{i=1}^{k}u_{i}P_{i}\|_{2}$ and $\sum_{i=1}^{k}c_{i}=1$. Further details about the data encoding and the algorithmic flow of this variational quantum algorithm can be found in Appendix B.
@@
We benchmark the variational algorithm using Qiskit 1.3.1 [121] with a simulator (Qiskit Aer MPS) and actual quantum hardware (ibm__\__cusco). All classical execution steps were run on a Macbook Pro (M1 Max CPU and 32 GB RAM). Each run of the algorithm consists of 10 iterations, where each iteration involves drawing 1024 shots from the quantum computer and hyperparameter optimization of the respective coefficients with a black-box solver, as explained above. We run the algorithm once for all instances with size $n=4$ and $n=5$, and present a performance summary in Figures 19. Recall from Section 4.3.3 that there are 10 instances per matrix size and density (sparse/dense) and that sparse/dense matrices are generated by sampling, respectively, $n$ and $n^{2}$ permutations matrices and weights uniformly at random. Furthermore, Table 13 presents an example of a submission for a particular problem instance (B5_5_8) where we run the algorithm 10 times.
@@
Figure 19 shows the simulation and hardware results for $n=4$ and $n=5$ with sparse and dense matrices; the dots indicate the number of permutations in the decomposition (left y-axis), and the squares/crosses represent the decomposition error (right y-axis). The x-axis indicates the problem instance BX_Y_Z, where X is the size of the doubly stochastic matrix (i.e., $n$), Y the number of permutations used to generate the doubly stochastic matrix ($n$ for sparse and $n^{2}$ for dense), and Z the instance number. Observe from both figures that when $n=4$ and the target matrix is sparse, the variational algorithm can consistently obtain sparse decompositions of length 3 and 4. For both simulation and hardware results, the solutions coincide with the solution obtained with CPLEX except for one instance (B4_4_9; simulation), where the decomposition length is 5 instead of 4 (CPLEX). Also, the decomposition is exact for all instances (i.e., zero error). With $n=4$ and a dense target matrix, the performance remains similar to the sparse case. Finally, with $n=5$, the performance degrades significantly compared to the $n=4$ case, with only a few instances reaching sparse decompositions with decomposition errors below $10^{-10}$. Despite solving only a few instances, the result is remarkable for a variational algorithm since there are $\binom{5!}{5}\sim 190$ million combinations of 5 permutation matrices of size $5$. Moreover, some of the decompositions obtained with the variational algorithm are better than the decompositions obtained with approximate classical algorithms such as Birkhoff+ and Blended FW. For instance, with instance B5_5_6, Birkhoff+ obtains an exact decomposition of length 7, while the variational algorithm obtains an exact decomposition of length 5.
Problem Identifier B5__\__5__\__8 ($n=5$; density: sparse; instance number: 8)
@@
Decision Variables Range $[0,10000]$ (black-box step 1); $[0,1]$ (black-box step 2)
@@
Workflow Eeach iteration of the variational algorithm calls:
@@
Table 13: (Birkhoff) Report of benchmark submission for instance B5_5_8 ($n=5$; density: sparse; instance number: 8) from the Minimum Birkhoff Decomposition problem collection provided in QOBLIB [193]. All runtimes are reported as the average over the 10 different runs.
@@ 5.3 Maximum Independent Set
We neglect the irrelevant constant energy offsets, i.e., terms that are neither linear nor quadratic in $Z_{i}$. These terms are $|V|/2$, and $-\lambda|E|/4$ and stem from the objective and constraints, respectively. Therefore, the total cost function Hamiltonian is $H=H_{\text{obj}}+\lambda H_{\text{const}}$. To produce bitstrings from the quantum hardware, we sample from a QAOA trial wavefunction $|{\psi(\beta,\gamma)}\rangle=e^{-i\beta H_{M}}e^{-i\gamma H^{\prime}}|+\rangle$. Here, $H_{M}=\sum_{i\in V}X_{i}$ is the standard QAOA mixer and $H^{\prime}$ is the cost operator $H_{\text{obj}}+\lambda H_{\text{const}}^{\prime}$ used to design the trial wavefunction and we apply $p=1$ repetitions. Crucially, $H_{\text{const}}^{\prime}$ may implement only a subset of the constraints in $H_{\text{const}}$ to limit the depth of the quantum circuit. We must now optimize the parameters $\beta$, $\gamma$, and $\lambda$. In a typical QAOA, this is done in a closed loop with the quantum hardware, which is time and resource-consuming. Best practices on how to carefully choose transpilation and error suppression strategies for a QAOA benchmark execution can be found on Github [20]. In our case, we choose to optimize $\beta$, $\gamma$, and $\lambda$ on classical hardware by evaluating $\langle H\rangle$, which is efficiently possible for a depth-one ansatz [69]. Optimizing $\lambda$ is a hard task [8]. Here, we chose to strike a balance between the impact of the actual objective part of the Hamiltonian and the constraint part of the Hamiltonian. This can be formulated, e.g., as the optimization problem
@@
$\max_{\beta,\gamma}\leavevmode\nobreak\ \langle\psi(\beta,\gamma)|H_{\text{obj}}+\lambda H_{\text{const}}|\psi(\beta,\gamma)\rangle=\max_{\beta,\gamma}\leavevmode\nobreak\ E(\beta,\gamma;\lambda).$ (33)
@@
We would require $\lambda>1$ to make the lowest energy state of $H$ correspond to the optimal solution of the MIS problem. However, we can only afford large values of lambda if we run deep QAOAs with $p>1$ since they have a better resolution than QAOAs with $p=1$. Since hardware noise restricts us to shallow circuits, we use $p=1$ and work with small values of $\lambda$. Once the optimal parameters are found we create a quantum circuit corresponding to $|\psi(\beta^{\star}(\lambda^{\star}),\gamma^{\star}(\lambda^{\star});\lambda^{\star})\rangle$ and sample 1024 candidate solutions on ibm_fez, a 156 qubit Heron R2 superconducting qubit processor [116]. These circuits are executed in session mode since this allows us to report the QPU time as the payload quantum circuit preparation and execution time–as discussed in Section 3.3.2. Finally, the samples are post-processed by the greedy classical method described in Appendix C. This benchmark was executed with Numpy 1.23.5, Qiskit 1.2.2, and SciPy 1.11.3 running on Python 3.11.5. We benchmark this approach on stable set instances referred to as mamila kangoroo and aves sparrow with 17 and 52 variables, which are given in QOBLIB [193]. All constraints of the 17-variable instance are implemented on hardware, i.e., $H^{\prime}_{\text{const}}=H_{\text{const}}$. This requires a total of 308 CNOT gates. For the 52-variable instance, we only implement 16.7% of the constraints by allowing three layers of SWAP gates [245]. We, therefore, only implement the constraints that can be built up on a line of qubits on which we apply three layers of SWAP gates as described in Ref. [245]. The resulting circuit has a total of 207 CNOT gates. This keeps the fidelity of the circuit, approximated by $(1-E_{\text{CZ}})^{n_{\text{CZ}}}$, above 50%. Here, $E_{\text{CZ}}$ and $n_{\text{CZ}}$ are the median error per two-qubit CZ gate and the number of two-qubit CZ gates, respectively. With the 52-variable instance, we do not sample any feasible solutions. However, after the greedy classical post-processing, we obtained the optimal sample. The runtime results from the process described above are reported in Table 14. As expected, most of the runtime is dedicated to optimizing the parameters $\beta$, $\gamma$, and $\lambda$. We observe that 60% and 27% of the post-processed samples for the 17 and 52 decision variable problems, respectively, are optimal, see Figure 20. Finally, Table 15 presents an example of a submission report for the $52-$variable instance.
@@
    Lagrange multiplier are optimized classically.
@@
Table 15: (Independent Set) Benchmark submission, for instance aves-sparrow-social from the independent set benchmarking problem collection provided in the QOBLIB repository [193].

### v2

@@ 5.1 Low Autocorrelation Binary Sequences
Figure 17: (LABS) Comparison of the time-to-solution for LABS instance ranging from size 10 to 30 using BF-DCQO and QAOA across different numbers of layers: $p=1$, $p=4$, $p=8$, and $p=12$. The data for QAOA is sourced from [219]. Note that the BF-DCQO times reflect idealized QPU runtime estimates, excluding transpilation.
The quantum adiabatic algorithm (QAA) evolves an easily prepared ground state of an initial Hamiltonian $H_{i}$ to the unknown ground state of a target Hamiltonian $H_{f}$, which should correspond to the optimal solution of an optimization problem of interest. This evolution is described by a time-dependent Hamiltonian of the form
@@
An extension of DCQO is bias-field DCQO (BF-DCQO), which is an iterative algorithm [44, 211, 123]. In BF-DCQO, the information from the final evolved state is fed back as a bias for a new initial state, such that the initial Hamiltonian for another DCQO iteration is adapted accordingly. In addition to that, we consider a classical post-processing scheme where we perform a local search [60, 67, 254] on the $n_{keep}$ lowest energy bitstrings obtained from BF-DCQO at each iteration. This local search can be seen as a zero-temperature simulated annealing with $n_{s}$ sweeps.
@@
Figure 18: (LABS) Comparison of the time-to-solution for BF-DCQO, Gurobi, and CPLEX. The data for the classical optimizers was sourced from [219]. Note that the BF-DCQO times reflect idealized QPU runtime estimates, excluding transpilation.
@@
Table 11: (LABS) Summary of results for ten experimental trials of BF-DCQO on the size $n=20$ LABS problem executed on ibm_\_marrakesh. The overall runtime consists of pre-processing time, which includes circuit construction and transpilation, QPU runtime, which includes the payload circuit preparation and execution time, and post-processing time, which includes the local search post-processing. Further $p_{gs}$ shows the probability of sampling the best solution–here the optimal one.
To test the performance experimentally, we selected the LABS problem with a sequence length of $n=20$, which exceeds the largest LABS problem implemented on a quantum computer to date ($\text{i.e.}\;n=18$ [219]), and present the results in the submission format introduced in Table 10 in Table 12.
@@
Table 12: (LABS) Report of benchmark submission for the LABS instance with $n=20$, also available in [202]. All runtimes are reported as the average over the 10 runs.
Since the Hamiltonian corresponding to LABS is dense with both 2-local and 4-local terms, its realization is challenging due to the large number of entangling gates required. Therefore, we restrict the counterdiabatic terms to include only 2-local terms. We used the 156-qubit ibm_\_marrakesh device with optimization level 2 transpilation, dynamical decoupling based on an XpXm sequence, and Qiskit 1.3.1. To facilitate the experimental implementation, we adopted the following strategy for initializing the bias fields. First, we created a random bitstring and applied a local greedy search with $n_{\text{s}}=5$. Then, we use the resulting bitstring to compute the initial bias fields. We set the initial number of shots to $n_{\text{shots}}=500$ and $n_{\text{s}}=2$. The $n_{\text{keep}}$ parameter is chosen as $50\%$ and $n_{\text{shots}}$ is increased by $500$ after each iteration. In Table 11, we present the results for ten different bias-field initializations. For the selected set of parameters, all the initializations resulted in a non-zero success probability. The $n=20$ LABS problem has an eight-fold degenerate ground state, which can be divided into two distinct independent groups based on spin-flip and reflection symmetry. In our results, we measure optimal sequences from both groups. The average overall runtime to measure the optimal solution was $61.2$ seconds. However, excluding the payload quantum circuit preparation, the average runtime becomes $9.1$ seconds, which is still better than the two exact classical methods shown in Figure 18.
@@ 5.2 Minimum Birkhoff Decomposition
Figure 19: (Birkhoff) Quantum simulation (top) hardware (bottom) results for $n=4$ (left) and $n=5$ (right). The error markers with $10^{-12}$ error indicate an exact decomposition (i.e., zero error). The x-axis indicates the problem instance BX_sparse/dense_Z, where X is the size of the doubly stochastic matrix (i.e., $n$), sparse/dense is the number of permutations used to generate the doubly stochastic matrix ($n$ for sparse and $n^{2}$ for dense), and Z is the instance number.
@@
In the experiments presented below, we employed Optuna 4.1 [10] as a black-box solver for Eq. (29) to optimize the parameters $\theta$. After sampling the permutation matrices, the coefficients $c_{1},\dots,c_{k}$ are computed classically in two steps. In the first step (black-box step 1), we use CPLEX 22.1.1 [51] to minimize $\|sD-\sum_{i=1}^{k}u_{i}P_{i}\|_{2}$ with respect to $u_{i}\in\{0,1,2,\dots,s\}$ subject to $0\leq u_{i}\leq s$ and $\sum_{i=1}^{k}u_{i}=s$ where $s$ is a given integer in the dataset that makes $sD$ and interger-valued matrix. In the second step (black-box step 2), we use again CPLEX 22.1.1 to find the coefficients $c_{i}\in[0,1]$ that provide the smallest decomposition subject to $\|D-\sum_{i=1}^{k}c_{i}P_{i}\|_{2}\leq\|D-\frac{1}{s}\sum_{i=1}^{k}u_{i}P_{i}\|_{2}$ and $\sum_{i=1}^{k}c_{i}=1$. Further details about the data encoding and the algorithmic flow of this variational quantum algorithm can be found in Appendix D.
@@
We benchmark the variational algorithm using Qiskit 1.3.1 [127] with a simulator (Qiskit Aer MPS) and actual quantum hardware (ibm_\_cusco). All classical execution steps were run on a MacBook Pro (M1 Max CPU and 32 GB RAM). Each run of the algorithm consists of 10 iterations, where each iteration involves drawing 1024 shots from the quantum computer and hyperparameter optimization of the respective coefficients with a black-box solver, as explained above. We run the algorithm once for all instances with size $n=4$ and $n=5$, and present a performance summary in Figures 19. Recall from Section 4.3.3 that there are 10 instances per matrix size and density (sparse/dense) and that sparse/dense matrices are generated by sampling, respectively, $n$ and $n^{2}$ permutation matrices and weights uniformly at random. Furthermore, Table 13 presents an example of a submission for a particular problem instance (B5_5_8) where we run the algorithm 10 times.
@@
Figure 19 shows the simulation and hardware results for $n=4$ and $n=5$ with sparse and dense matrices; the dots indicate the number of permutations in the decomposition (left y-axis), and the squares/crosses represent the decomposition error (right y-axis). The x-axis indicates the problem instance BX_Y_Z, where X is the size of the doubly stochastic matrix (i.e., $n$), Y is the number of permutations used to generate the doubly stochastic matrix ($n$ for sparse and $n^{2}$ for dense), and Z is the instance number. Observe from both figures that when $n=4$ and the target matrix is sparse, the variational algorithm can consistently obtain sparse decompositions of length 3 and 4. For both simulation and hardware results, the solutions coincide with the solution obtained with CPLEX except for one instance (B4_4_9; simulation), where the decomposition length is 5 instead of 4 (CPLEX). Also, the decomposition is exact for all instances (i.e., zero error). With $n=4$ and a dense target matrix, the performance remains similar to the sparse case. Finally, with $n=5$, the performance degrades significantly compared to the $n=4$ case, with only a few instances reaching sparse decompositions with decomposition errors below $10^{-10}$. Despite solving only a few instances, the result is remarkable for a variational algorithm since there are $\binom{5!}{5}\sim 190$ million combinations of 5 permutation matrices of size $5$. Moreover, some of the decompositions obtained with the variational algorithm are better than the decompositions obtained with approximate classical algorithms such as Birkhoff+ and Blended FW. For instance, with instance B5_5_6, Birkhoff+ obtains an exact decomposition of length 7, while the variational algorithm obtains an exact decomposition of length 5.
Problem Identifier B5_\_5_\_8 ($n=5$; density: sparse; instance number: 8)
@@
Workflow Each iteration of the variational algorithm calls:
@@
Table 13: (Birkhoff) Report of benchmark submission for instance B5_5_8 ($n=5$; density: sparse; instance number: 8) from the Minimum Birkhoff Decomposition problem collection provided in [202]. All runtimes are reported as the average over the 10 different runs.
@@ 5.3 Maximum Independent Set
We neglect the irrelevant constant energy offsets, i.e., terms that are neither linear nor quadratic in $Z_{i}$. These terms are $|V|/2$, and $\lambda|E|/4$ and stem from the objective and constraints, respectively. Therefore, the total cost function Hamiltonian is $H=H_{\text{obj}}+\lambda H_{\text{const}}$. To produce bitstrings from the quantum hardware, we sample from a QAOA trial wavefunction $|{\psi(\beta,\gamma)}\rangle=e^{-i\beta H_{M}}e^{-i\gamma H^{\prime}}|+\rangle$. Here, $H_{M}=\sum_{i\in V}X_{i}$ is the standard QAOA mixer. $H^{\prime}$ is the cost operator $H_{\text{obj}}+\lambda H_{\text{const}}^{\prime}$ used to design the trial wavefunction, and we apply $p=1$ repetitions. Crucially, $H_{\text{const}}^{\prime}$ may implement only a subset of the constraints in $H_{\text{const}}$ to limit the depth of the quantum circuit. We must now optimize the parameters $\beta$, $\gamma$, and $\lambda$. In a typical QAOA, this is done in a closed loop with the quantum hardware, which is time and resource-consuming. Best practices on how to carefully choose transpilation and error suppression strategies for a QAOA benchmark execution can be found on Github [21]. In our case, we choose to optimize $\beta$, $\gamma$, and $\lambda$ on classical hardware by evaluating $\langle H\rangle$, which is efficiently possible for a depth-one ansatz [74]. Optimizing $\lambda$ is a hard task [8]. Here, we chose to strike a balance between the impact of the actual objective part of the Hamiltonian and the constraint part of the Hamiltonian. This can be formulated, e.g., as the optimization problem
@@
[…]
@@
We would require $\lambda>1$ to make the lowest energy state of $H$ correspond to the optimal solution of the MIS problem. However, we can only afford large values of lambda if we run deep QAOAs with $p>1$ since they have a better resolution than QAOAs with $p=1$. Since hardware noise restricts us to shallow circuits, we use $p=1$ and work with small values of $\lambda$. Once the optimal parameters are found, we create a quantum circuit corresponding to $|\psi(\beta^{\star}(\lambda^{\star}),\gamma^{\star}(\lambda^{\star});\lambda^{\star})\rangle$ and sample 1024 candidate solutions on ibm_fez, a 156 qubit Heron R2 superconducting qubit processor [122]. These circuits are executed in session mode as this allows us to report the QPU time as the payload quantum circuit preparation and execution time–as discussed in Section 3.3.2. Finally, the samples are post-processed by the greedy classical method described in Appendix E. This benchmark was executed using Numpy 1.23.5, Qiskit 1.2.2, and SciPy 1.11.3, all running on Python 3.11.5. We benchmark this approach on stable set instances referred to as mamila kangoroo and aves sparrow with 17 and 52 variables, which are given in [202]. All constraints of the 17-variable instance are implemented on hardware, i.e., $H^{\prime}_{\text{const}}=H_{\text{const}}$. This requires a total of 308 CNOT gates. For the 52-variable instance, we only implement 16.7% of the constraints by allowing three layers of SWAP gates [255]. We, therefore, only implement the constraints that can be built up on a line of qubits on which we apply three layers of SWAP gates as described in Ref. [255]. The resulting circuit has a total of 207 CNOT gates. This keeps the fidelity of the circuit, approximated by $(1-E_{\text{CZ}})^{n_{\text{CZ}}}$, above 50%. Here, $E_{\text{CZ}}$ and $n_{\text{CZ}}$ are the median error per two-qubit CZ gate and the number of two-qubit CZ gates, respectively. With the 52-variable instance, we do not sample any feasible solutions. However, after the greedy classical post-processing, we obtained the optimal sample. The runtime results from the process described above are reported in Table 14. As expected, most of the runtime is dedicated to optimizing the parameters $\beta$, $\gamma$, and $\lambda$. We observe that 60% and 27% of the post-processed samples for the 17 and 52 decision variable problems, respectively, are optimal, see Figure 20. Finally, Table 15 presents an example of a submission report for the $52$-variable instance.
@@
   Lagrange multipliers are optimized classically.
@@
Table 15: (Independent Set) Benchmark submission, for instance aves-sparrow-social from the independent set benchmarking problem collection provided in the [202] repository.